OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = x + 4 and passes through the point (−6, 3).

OpenStudy (anonymous):

Use the same method: Slope of the normal is -1 therefore \[y=-x-3\]

OpenStudy (anonymous):

y = −2/3x − 1

OpenStudy (anonymous):

the slope doesn't seem right to me

OpenStudy (anonymous):

y = 3/2x − 1

OpenStudy (anonymous):

y=-x-3

OpenStudy (anonymous):

y = 3/2x + 3